### v1

# Momentum and spin dynamics of photon in linear dielectric medium

Abstract
The momentum of light in dielectric media has been a century-long controversy that continues to attract significant interest. In a linear dielectric medium with refractive index $n$, the momentum is predicted to be smaller by a factor of $n$ according to Max Abraham, and larger by the same factor according to Hermann Minkowski. By studying the coupled dynamics of photons and dipoles in a dielectric medium, we show that the change in momentum of the dipole corresponds to the Abraham momentum. On the other hand, the Minkowski momentum determines the direction of propagation and the precession of the photon spin, which is described by an SO(3) rotation.
Introduction. Due to the opposing predictions by Max Abraham and Hermann Minkowski, the momentum of light in dielectric media has been a century-long debate [1, 2, 3]. Since the wavelength in a dielectric medium with refractive index $n$ is shortened, Minkowski predicted that the wavevector $k=2\pi n/\lambda$ and momentum are larger:
$p_{\mathrm{M}}=\hbar k=\frac{nhc}{\lambda}.$ (1)
On the other hand, Abraham argued that the reduction in velocity $v=c/n$ should reduce the momentum:
$p_{\mathrm{A}}=\frac{\hbar\omega}{c^{2}}v=\frac{hc}{n\lambda}.$ (2)
While some experiments support one prediction [4, 5, 6, 7, 8, 9, 10, 11], no particular setup can conclusively disprove the other [12, 13, 14].
Theoretical discussions of the Abraham-Minkowski controversy often employ field theory approaches [15, 16], which are well suited for describing the wave properties of particles. These discussions tend to focus on momentum density [17, 18, 19]. However, to properly study the momentum of light, it is necessary to consider the particle nature of the wave, e.g., photon-electron elastic collisions in Compton scattering. In Compton scattering, the momentum of light is clearly determined from the change in electron momentum [20].
In this article, we focus on a linear dielectric medium modeled as a Drude material, where electrons are connected by springs. This model yields a linear refractive index for photons at low frequencies [21]. The Drude model has been successfully used to describe electric, heat, and optical transport properties in materials [22]. The theoretical predictions of the Drude model are in reasonable agreement with observed values for many materials [23] and have been applied to describe Hall effects [24, 25].
Refractive index in linear dielectric medium. The classical Drude model focuses on the dynamics of electrons under an electric field $\mathbf{E}$. Here, we include the dynamics of the medium by modeling the positive charge as connected to the electron by a spring with natural oscillation frequency $\omega_{0}$, yielding coupled equations:
$m_{\pm}\frac{d^{2}\mathbf{r}_{\pm}}{dt^{2}}=-m\omega_{0}^{2}\left(\mathbf{r}_{\pm}-\mathbf{r}_{\mp}\right)\pm e\mathbf{E}\pm e\frac{d\mathbf{r}_{\pm}}{dt}\times\mathbf{B},$ (3)
where $m_{-}$ and $m_{+}\gg m_{-}$ are the masses of negative and positive charges, respectively, and $m=(m_{+}^{-1}+m_{-}^{-1})^{-1}$ is the effective mass of the relative dipole motion. The first term on the right-hand side represents the internal restoring force of the dipole, with natural oscillation frequency $\omega_{0}$. $\mathbf{B}$ is the magnetic field.
By defining the center of mass $\mathbf{R}=(m_{+}\mathbf{r}_{+}+m_{-}\mathbf{r}_{-})/M$ and relative position $\mathbf{r}_{e}=\mathbf{r}_{-}-\mathbf{r}_{+}$, one obtains
$M\frac{d^{2}\mathbf{R}}{dt^{2}}$ $=-e\frac{d\mathbf{r}_{e}}{dt}\times\mathbf{B},$ (4)
$\frac{d^{2}\mathbf{r}_{e}}{dt^{2}}$ $=-\omega_{0}^{2}\mathbf{r}_{e}-\frac{e}{m}\mathbf{E}-\frac{e}{m}\frac{d\mathbf{R}}{dt}\times\mathbf{B},$ (5)
where $M=m_{+}+m_{-}$. Equation 4 describes the center of mass motion influenced by the net Lorentz force on the dipole, and will be used to determine the photon momentum in the medium.
[…]
$\mathbf{r}_{e}$ $=\frac{-e}{m(\omega_{0}^{2}-\omega^{2})}\mathbf{E}.$ (6)
The relative permittivity, related to the electric polarization $\mathbf{P}=-e\mathbf{r}_{e}/V$, is then
$\varepsilon_{r}$ $=1+\frac{\partial\mathbf{P}}{\varepsilon_{0}\partial\mathbf{E}}=1+\frac{\omega_{p}^{2}}{\omega_{0}^{2}-\omega^{2}},$ (7)
where $\omega_{p}=\sqrt{e^{2}/(\varepsilon_{0}mV)}$ is the plasma frequency, $1/V$ is the density, and $\varepsilon_{0}$ is the vacuum permittivity. In the low-frequency regime, the system behaves as a linear medium with refractive index
$n=\lim_{\omega\ll\omega_{0}}\sqrt{\varepsilon_{r}}=\sqrt{1+\frac{\omega_{p}^{2}}{\omega_{0}^{2}}}.$ (8)
Equation of motion of photon. Coupled equations of dipole and photon arise from the Drude equation and Maxwell’s equations for current $\mathbf{j}=\dot{\mathbf{P}}=-e\dot{\mathbf{r}}_{e}/V$:
$\frac{\partial}{\partial t}\left[\begin{array}[]{c}\mathbf{E}\\ c\mathbf{B}\\ \dot{\mathbf{r}}_{e}\\ \mathbf{r}_{e}\end{array}\right]=\left[\begin{array}[]{cccc}0&\frac{ic}{\hbar}\hat{\mathbf{p}}\cdot\mathbf{s}&\frac{e}{\varepsilon_{0}V}&0\\ -\frac{ic}{\hbar}\hat{\mathbf{p}}\cdot\mathbf{s}&0&0&0\\ -\frac{e}{m}&0&0&-\omega_{0}^{2}\\ 0&0&1&0\end{array}\right]\left[\begin{array}[]{c}\mathbf{E}\\ c\mathbf{B}\\ \dot{\mathbf{r}}_{e}\\ \mathbf{r}_{e}\end{array}\right].$ (9)
Here, $\hat{\mathbf{p}}=-i\hbar\nabla$, and $\mathbf{s}$ is a Hermitian SO(3) rotation representation [26] fulfilling the angular momentum commutation relations $[s_{i},s_{j}]=i\sum_{k}\epsilon_{ijk}s_{k}$ for spin-1, with $s^{2}=2$, and $s_{z}=0,\pm 1$.
Equation 9 can be written in Schrödinger form:
$i\hbar\frac{\partial}{\partial t}\Psi=\hat{H}\Psi=\left[\begin{array}[]{cc}\hat{H}_{c}&\hat{H}_{ac}\\ \hat{H}_{ac}^{t}&\hat{H}_{a}\end{array}\right]\Psi,$ (12)
where
$\Psi=\left[\begin{array}[]{cccc}\mathbf{c}_{+}&\mathbf{c}_{-}&\mathbf{a}&\mathbf{a}^{\dagger}\end{array}\right]^{t},$
with $\mathbf{c}_{\pm}=\sqrt{\frac{\varepsilon_{0}V}{2\hbar\omega_{0}}}\left(\mathbf{E}\pm ic\mathbf{B}\right)$ representing the right-handed and left-handed photon wave functions [27], and
$\mathbf{a}=\sqrt{\frac{m\omega_{0}}{2\hbar}}\left(\mathbf{r}_{e}+i\frac{\dot{\mathbf{r}}_{e}}{\omega_{0}}\right),\quad\mathbf{a}^{\dagger}=\sqrt{\frac{m\omega_{0}}{2\hbar}}\left(\mathbf{r}_{e}-i\frac{\dot{\mathbf{r}}_{e}}{\omega_{0}}\right),$
analogous to annihilation and creation operators of a quantum harmonic oscillator.
The uncoupled Hamiltonians for the photon and dipole are
$\hat{H}_{c}$ $=\left[\begin{array}[]{cc}c\hat{\mathbf{p}}\cdot\mathbf{s}&0\\ 0&-c\hat{\mathbf{p}}\cdot\mathbf{s}\end{array}\right],\quad\hat{H}_{a}=\left[\begin{array}[]{cc}\hbar\omega_{0}&0\\ 0&-\hbar\omega_{0}\end{array}\right],$ (17)
and the coupling Hamiltonian is
$\hat{H}_{ac}=\frac{\hbar\omega_{p}}{2}\left[\begin{array}[]{cc}1&-1\\ 1&-1\end{array}\right].$ (20)
$\hat{H}_{ac}$ enables indirect coupling between right- and left-handed photons. Since $\hat{H}$ is Hermitian, the Heisenberg equation can be used to find time derivatives: $df/dt=[f,\hat{H}]/(i\hbar)$.
Figure 1: Minkowski diagram illustrating the Zitterbewegung of a photon in a linear dielectric medium due to indirect coupling between right-handed ($\mathbf{c}_{+}$) and left-handed ($\mathbf{c}_{-}$) photons mediated by dipole oscillation ($\mathbf{a}$).
The photon velocity operator is obtained as the momentum derivative of the Hamiltonian,
$\hat{\mathbf{v}}=\frac{\partial\hat{H}}{\partial\mathbf{p}}=\left[\begin{array}[]{cccc}c\mathbf{s}&0&0&0\\ 0&-c\mathbf{s}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right].$ (25)
The spin dynamics satisfy
$\frac{d\mathbf{s}}{dt}=\frac{1}{i\hbar}[\mathbf{s},\hat{H}]=\frac{1}{\hbar}\hat{\mathbf{p}}\times\hat{\mathbf{v}},$ (26)
ensuring conservation of total angular momentum $d(\hbar\mathbf{s}+\hat{\mathbf{r}}\times\hat{\mathbf{p}})/dt=0$. Equation 26 implies $d(\hat{\mathbf{p}}\cdot\mathbf{s})/dt=0$, so $s_{z}$ is a good quantum number when $\hat{z}\parallel\hat{\mathbf{p}}$. Hence, the eigenvalue of spin along $\hat{z}\parallel\hat{\mathbf{p}}$ determines the observed photon velocity: right-handed (left-handed) photon velocity is parallel (anti-parallel) to $s_{z}$. The superposition of right- and left-handed photons, indirectly coupled via dipole oscillations, determines the photon’s velocity in the dielectric medium, as shown in Fig. 1:
$\langle\hat{\mathbf{v}}\rangle=\Psi^{\dagger}\hat{\mathbf{v}}\Psi=c\left(\mathbf{c}_{+}^{\dagger}\mathbf{s}\mathbf{c}_{+}-\mathbf{c}_{-}^{\dagger}\mathbf{s}\mathbf{c}_{-}\right).$ (27)
This phenomenon resembles the zitterbewegung of electrons governed by the Dirac equation [28], except that the coupling between right- and left-handed photons here is mediated indirectly by dipole oscillations.
Minkowski momentum in refraction. For small momenta $pc\ll\hbar\omega_{0}$, the Hamiltonian can be diagonalized as
$\hat{H}=U\left[\begin{array}[]{cccc}\frac{c}{n}\mathbf{s}\cdot\hat{\mathbf{p}}&0&0&0\\ 0&-\frac{c}{n}\mathbf{s}\cdot\hat{\mathbf{p}}&0&0\\ 0&0&n\hbar\omega_{0}&0\\ 0&0&0&-n\hbar\omega_{0}\end{array}\right]U^{t},$ (32)
with
$U=\frac{\left[\begin{array}[]{cccc}1+n&1-n&-\sqrt{n^{2}-1}&-\sqrt{n^{2}-1}\\ 1-n&1+n&-\sqrt{n^{2}-1}&-\sqrt{n^{2}-1}\\ -\sqrt{n^{2}-1}&-\sqrt{n^{2}-1}&-n-1&n-1\\ -\sqrt{n^{2}-1}&-\sqrt{n^{2}-1}&n-1&-n-1\end{array}\right]}{2n}.$
Here, $n$ is the refractive index from Eq. 8. The dispersion relations, corresponding to eigenfrequencies of the coupled system, are shown in Fig. 2. The first eigenvalue corresponds to the photon dispersion in the linear medium, yielding the Minkowski momentum $\hat{p}=\frac{n\hbar\omega}{c}$. The change in the momentum operator during refraction from vacuum to the linear dielectric medium conserves the tangential momentum $\hat{\mathbf{p}}_{t}$ at the refraction surface, since
$\frac{d\mathbf{p}}{dt}=-\nabla\hat{H}.$
Figure 2: Dispersion in a linear dielectric medium: hybridization of right-handed (blue dashed line), left-handed photons (red dashed line), and dipole (green dotted line) modes results in linear dispersion $\hbar\omega=\pm pc/n$ (purple lines) at low frequencies.
The eigenvector $U_{1}$ associated with the first eigenvalue determines the contributions of $\mathbf{c}_{+}$ and $\mathbf{c}_{-}$ in the observed photon velocity:
$\langle\hat{\mathbf{v}}\rangle_{1}=U_{1}^{t}\hat{\mathbf{v}}U_{1}=\mathbf{s}\frac{c}{n}.$
The relation between velocity and spin governs spin precession:
$\frac{d\mathbf{s}}{dt}=\hat{\mathbf{p}}\times\langle\hat{\mathbf{v}}\rangle_{1}=\frac{c}{n\hbar}\hat{\mathbf{p}}\times\mathbf{s}.$ (33)
The spin eigenvalue $s_{z}=1$ along $\hat{z}\parallel\hat{\mathbf{p}}$ determines the photon velocity $c/n$ in the medium. In metamaterials with negative refractive index, the left-handed photon contribution dominates, reversing the direction of velocity relative to $\mathbf{s}$ and $\hat{\mathbf{p}}$ (Fig. 3).
Figure 3: (a) Refraction conserves tangential momentum $\hat{\mathbf{p}}_{t}$ at the interface. The photon velocity $\hat{\mathbf{v}}$ in the medium is a superposition of right-handed ($\mathbf{c}_{+}$) and left-handed ($\mathbf{c}_{-}$) photons, whose velocities are parallel and anti-parallel to the spin $\mathbf{s}$, respectively. The spin $\mathbf{s}$ precesses around $\hat{\mathbf{p}}$ with $s_{z}\parallel\hat{\mathbf{p}}$. (b) In metamaterials, the left-handed photon ($\mathbf{c}_{-}$) contribution dominates, resulting in velocity opposite to $\hat{\mathbf{p}}$.
The transverse spin components $s_{x}$ and $s_{y}$ determine the photon phase, as illustrated in Fig. 3. Since the momentum $\hat{p}$ in the dielectric medium is $n$ times larger than in vacuum, the precession frequency remains unchanged. However, because the speed is reduced by $n$, the wavelength of the photon in the medium is correspondingly shorter by the same factor.
Abraham momentum and acceleration. Thus far, we have identified conserved quantities of the photon in the medium: the momentum operator $\hat{p}$, the spin component $s_{z}$ parallel to it, and energy. The photon energy entering the medium $\hbar\omega$ equals to the sum of electromagnetic and dipole oscillation energies,
$\hbar\omega=\frac{V\varepsilon_{0}}{2}\left(E^{2}+c^{2}B^{2}\right)+\frac{1}{2}\left(m\omega_{0}^{2}r_{e}^{2}+m\dot{r}_{e}^{2}\right).$
This relation links acceleration and momentum change. The acceleration operator satisfies
$\frac{d\hat{\mathbf{v}}}{dt}=$ $\frac{c^{2}}{\hbar}\hat{\mathbf{p}}\times\left[\begin{array}[]{cccc}\mathbf{s}&0&0&0\\ 0&\mathbf{s}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right]+\frac{c\hbar\omega_{p}}{2}\left[\begin{array}[]{cccc}0&0&\mathbf{s}&-\mathbf{s}\\ 0&0&-\mathbf{s}&\mathbf{s}\\ -\mathbf{s}&\mathbf{s}&0&0\\ \mathbf{s}&-\mathbf{s}&0&0\end{array}\right],$ (42)
$\frac{d\langle\hat{\mathbf{v}}\rangle}{dt}$ $=\Psi^{\dagger}\frac{d\hat{\mathbf{v}}}{dt}\Psi=\frac{e\dot{\mathbf{r}}_{e}\times c\mathbf{B}}{\frac{V\varepsilon_{0}}{2}\left(E^{2}+c^{2}B^{2}\right)+\frac{1}{2}\left(m\omega_{0}^{2}r_{e}^{2}+m\dot{r}_{e}^{2}\right)}.$ (43)
This relates acceleration to the reaction force on the dipole Lorentz force in Eq. 4:
$\frac{d}{dt}\frac{\hbar\omega}{c^{2}}\langle\hat{\mathbf{v}}\rangle=e\dot{\mathbf{r}}_{e}\times\mathbf{B}.$ (44)
We conclude that the quantity $\hbar\omega\langle\hat{\mathbf{v}}\rangle/c^{2}$ governs momentum transfer to the dipole. Its magnitude matches the Abraham momentum of a photon in a linear medium:
$\frac{\hbar\omega}{c^{2}}\langle\hat{\mathbf{v}}_{1}\rangle=\frac{\hbar\omega}{nc^{2}}s_{z}\hat{z}=\frac{\hbar\omega}{nc}\hat{z}.$
Conclusion. In summary, we have analyzed the photon dynamics in a dielectric medium through the coupled Hamiltonian of photons and dipoles. The dipole oscillation in the medium mediates an indirect coupling between right-handed and left-handed photons, whose velocities are parallel and antiparallel to the spin $\mathbf{s}$, respectively.
Conservation of spin ensures that the superposition of these photons reduces the photon velocity in the medium, with $s_{z}$ aligned parallel to the momentum operator $\hat{\mathbf{p}}$. The precession of transverse spin components $s_{x,y}$ around $\hat{\mathbf{p}}$ determines the photon phase. Refraction into the dielectric medium conserves the precession frequency and the tangential momentum $\hat{\mathbf{p}}_{t}$, thereby increasing the magnitude of $\hat{\mathbf{p}}$. This momentum corresponds to the Minkowski momentum, which sets the direction of spin and thus the direction of velocity.
Furthermore, the Abraham momentum governs the actual transfer of momentum to the medium, as the acceleration of the photon (Eq. 44) is directly related to the Lorentz force reaction on the dipole (Eq. 4). Our results clarify the coexistence of both Abraham and Minkowski momenta in dielectric media, assigning distinct physical roles to each.